write the distiction between alpha, beta and gamma rays taking into accounnt its charge , mass , identity and penetrating power...

Grade:11

1 Answers

Aarti Gupta
8 years ago
An unstable nucleus or an radioactive one decays to give three types of radiations which are – alpha,beta and gamma rays.
Their properties are as follows--
1). Alpha rays – These are helium ions and represented as 2He4.Charge = +2 and mass = 4.Their penetrating power is less but very high ionising power.
2).Beta rays -- Thes are high kinetic energy electrons which are represented as 0e-1.Mass = 1/1850 and charge = --1.These rays have moderate penetration and moderate ionsing power.
3).Gamma rays –These are very high frequency electromagnetic radiation and represented as 0\gamma0 .Mass as well as charge of these radiations are 0.Very high penetrating power and lowest ionising power among all three radiations.
When these radiations hit living cells they cause ionisation effects,can kill cells directly and cause genetic damage.